OpenStudy (anonymous):

Find the indefinite integral of 1/xln(x^3)

OpenStudy (anonymous):

let t=ln(x^3) \[\Large dt=\frac{3x^2}{x^3}dx\] \[\Large dt=\frac{3}{x}dx\] \[\Large \frac{1}{3}dt=\frac{1}{x}dx\] \[\Large \int\limits_{}^{} \frac{1}{3t}dt=\frac{1}{3}\ln(t)+c\] now substitute bacl t=ln(x)^3) to get the answer .

OpenStudy (anonymous):

Oh! Thank you so much!